### v2

@@ Abstract
Atypical mitosis marks a deviation in the cell division process that has been shown be an independent prognostic marker for tumor malignancy. However, atypical mitosis classification remains challenging due to low prevalence, at times subtle morphological differences from normal mitotic figures, low inter-rater agreement among pathologists, and class imbalance in datasets. Building on the Atypical Mitosis dataset for Breast Cancer (AMi-Br), this study presents a comprehensive benchmark comparing deep learning approaches for automated atypical mitotic figure (AMF) classification, including end-to-end trained deep learning models, foundation models with linear probing, and foundation models fine-tuned with low-rank adaptation (LoRA). For rigorous evaluation, we further introduce two new held-out AMF datasets - AtNorM-Br, a dataset of mitotic figures from the TCGA breast cancer cohort, and AtNorM-MD, a multi-domain dataset of mitotic figures from a subset of the MIDOG++ training set. We found average balanced accuracy values of up to 0.8135, 0.7788, and 0.7723 on the in-domain AMi-Br and the out-of-domain AtNorm-Br and AtNorM-MD datasets, respectively. Our work shows that atypical mitotic figure classification, while being a challenging problem, can be effectively addressed through the use of recent advances in transfer learning and model fine-tuning techniques. We make all code and data used in this paper available in this github repository: https://github.com/DeepMicroscopy/AMi-Br_Benchmark.
@@ 1 Introduction
Mitosis is the process where a cell replicates its genetic material (DNA) and then divides into two identical daughter cells. This process enables the proliferation of cells, supporting physiological tissue growth and facilitating the replacement of old, damaged, or worn-out cells. The dividing cell may be observed microscopically as an mitotic figure (MF) with specific morphologies resembling the highly regulated cell division phases. However, cell proliferation does not occur only under physiological conditions; it is also a hallmark of cancer, where cell division is dysregulated and increased. Subsequently, quantifying cell proliferation is essential for estimating the aggressiveness of many human and animal cancers, including human breast cancer (Fitzgibbons and Connolly, 2023; Kiupel et al., 2011; Louis et al., 2016). The routine method in the assessment of the growth rate is the mitotic count, which reflects the number of MFs within a specific tumor area. Elevated mitotic counts are associated with an increased risk of distant metastasis and higher risk of death (Medri et al., 2003).
@@
At the time of this writing, there have been very few studies investigating deep learning for AMFs (Aubreville et al., 2023a; Fick et al., 2024; Bertram et al., 2025). The automated classification of atypical vs. normal MFs is complicated by the substantial variation of morphologies, often with some mitotic figures displaying overlapping features of both normal and AMFs (Donovan et al., 2021; Bertram et al., 2023), and the low frequency of AMF, resulting in class imbalance. Given the prognostic relevance of differentiating normal vs. atypical MFs, standardized benchmarks, such as publicly available datasets, evaluation protocols, and baseline models, may serve as facilitators to push further developments in this field that tackle these challenges. In this context, foundation models offer a promising direction. These models, typically vision transformer (ViT)-based and trained on millions of histopathology slides using self supervised learning, have shown strong performance across various medical imaging tasks, including classification (Campanella et al., 2025; Breen et al., 2025), especially in small data regimes. These models are typically employed either via linear probing, where a lightweight classifier is trained on top of a frozen backbone, or through parameter-efficient fine-tuning techniques like Low Rank Adaptation (LoRA) (Hu et al., 2022). However, comprehensive studies leveraging these techniques for atypical vs. normal mitosis classification are still lacking.
The contribution of this work is two-fold: First, we present three datasets, designed for the automated identification of AMFs to establish suitable datasets that reflect the unique challenges of this task. Second, we provide a comprehensive benchmark comparing end-to-end trained deep learning architectures, foundation models with linear probing, and foundation models fine-tuned with LoRA for the task of atypical versus normal mitotic figure classification.
@@ 2 Datasets
Being the largest of the three datasets used in this study, we use data from the AMi-Br dataset for training in all evaluations. This allows us to use the other two newly introduced datasets as independent held-out sets. It has been shown that data distribution shifts (domain shifts) impede the performance of deep learning models significantly (Stacke et al., 2020; Aubreville et al., 2021). With these datasets, we can expect varying degrees of domain shifts, and which allows us to investigate generalization performance. Only a part of the AMi-Br dataset is used for training the models for binary classification of atypical vs. normal mitoses, and the rest is used as another held-out test dataset. The dataset statistics, including total number of atypical and normal MFs in each, annotation type, source dataset(s), species involved etc. have been summarized in Table 1.
@@
Property AMi-Br AtNorM-MD AtNorM-Br
Total MFs 3,720 2,107 746
Atypical 832 219 128
Normal 2,888 1,888 618
AMF Rate (%) 22.4% 10.4% 17.2%
Annotation Type 3-expert vote 5-expert vote Single expert
@@
Source Dataset(s) TUPAC16, MIDOG21 MIDOG++ TCGA (BRCA)
Species Human Human + Canine Human
@@ 2.2 AtNorM-MD
The second dataset, Atypical and Normal Mitosis (AtNorM)-MultiDomain (MD) extends beyond human breast cancer and covers six domains, spanning both human and canine tumors. These include canine lung cancer, canine lymphoma, canine cutaneous mast cell tumor, human neuroendocrine tumor, canine soft tissue sarcoma, and human melanoma. It was sampled randomly from all but the human breast cancer domains of the publicly available MIDOG++ dataset (Aubreville et al., 2023c). It comprises 2,107 MFs from 70 patients, with 219 (10.4%) being atypical. Labeling was performed as a majority vote of five pathology experts, three of whom were board-certified. We found total agreement per object by all five experts in 1,466 (69.6%) of the cases.
2.3 AtNorM-Br
The third dataset, AtNorM-Br, also made available publicly within the scope of this work, contains 746 MF instances from 179 patients from the breast cancer (BRCA) cohort of the The Cancer Genome Atlas (TCGA) (Lingle et al., 2016). TCGA contains images from various sources and with partially mixed quality, stemming from differences in staining protocols, scanning equipment, and tissue preparation across institutions, and is thus also a valuable asset for the assessment of generalization. This dataset was annotated by a single expert with high experience in AMF classification, according to which 128 MFs were found to be atypical and the remaining 618 were classified as normal, yielding an AMF rate of 17.16%.
@@ 3 Methods
3.1 End-to-end-trained Deep Learning Models
To establish robust baselines for atypical vs. normal mitoses classification, we evaluate three widely-used deep learning architectures – EfficientNetV2 (Tan and Le, 2021), ViT (Dosovitskiy et al., 2020), and Swin Transformer (Liu et al., 2021). All of these models have demonstrated strong performance across various medical imaging datasets and tasks. EfficientNetV2 is a well-established convolutional neural network (CNN) known for its strong performance and efficiency in various classification tasks. It serves as a representative of traditional CNN-based approaches. In contrast, the ViT represents a major shift from convolutional architectures to transformer-based models. A ViT processes images as sequences of patches and uses self-attention mechanisms to model global relationships across the input. This global receptive field allows the ViT to capture long-range dependencies, which can be particularly useful for complex histopathological patterns. The Swin Transformer further advances the transformer-based approach by introducing a hierarchical architecture with shifted windows. This architecture enables the model to compute attention locally while gradually building up to global representations, providing a balance between computational efficiency and the ability to capture both local details and global structure. Swin Transformers have been particularly effective in dense prediction tasks and high-resolution image analysis, making it well-suited for mitotic figure classification (Liu et al., 2021). The selected model architectures were chosen to be of similar size in parameters. Each model was trained in an end-to-end manner and evaluated on three held-out test datasets: a random, patient-stratified split of the AMi-Br dataset, accounting for $\approx$ 22 % of the samples in the dataset (samples not used in training or validation), and the entire AtNorM-MD and AtNorM-Br datasets.
@@ 3.2 Foundation Models
Foundation models are models that are trained on large and diverse corpora of data, often using unsupervised techniques like self supervision. The goal of foundation models is to train feature extractors that generalize well and are easily adaptable to down-stream tasks. In the field of computational pathology, we have seen a very strong incline towards publicly available foundation models, calling for an investigation on the task of AMF classification. We compare eight state-of-the-art models – UNI (Chen et al., 2024), UNI2-h (Chen et al., 2024; MahmoodLab, 2025), Virchow (Vorontsov et al., 2024), Virchow2 (Zimmermann et al., 2024), Prov-Gigapath (Xu et al., 2024), H-Optimus-0 (Saillard et al., 2024), H-Optimus-1 (Bioptimus, 2025), and H0-mini (Filiot et al., 2025). An overview of selected details for each model such as training dataset size, size in parameters and model type for training is provided in Table 2. All selected foundation models are ViT-based and have been pre-trained using the DINOv2 algorithm (Oquab et al., 2023) on datasets comprising of hundreds of thousands to multiple millions of whole slide images.
@@ 4 Experimental Setup
We set up detailed experiments for the above cases. The image preprocessing for the binary classification experiments included resizing all patches to 224×224 pixels. Training augmentations consisted of random horizontal flip, rotation, color jitter, and random resized crop, followed by normalization with ImageNet statistics. Validation images were only resized and normalized. The best checkpoint for each fold was selected based on validation balanced accuracy, which was used as the primary metric because of the class imbalance between the atypical and normal MFs in the dataset. All experiments were performed using 5-fold cross validation, where stratification of the data was based on the slide (patient) to ensure that all patches from a given slide appeared in only one split (train or validation), thus preventing data leakage. To mitigate class imbalance, we employed weighted random sampling during training, with sample weights inversely proportional to class frequencies. All models were trained using the Adam optimizer with L2 regularization (learning rate: $5\times 10^{-5}$, weight decay: $1\times 10^{-5}$) and cross-entropy loss. We applied early stopping with a patience of 15 epochs based on validation balanced accuracy, and used learning rate scheduling on plateau (factor: 0.5, patience: 3, minimum LR: $1\times 10^{-7}$). Each model was trained for up to 100 epochs with a batch size of 8. The rest of the case-specific details are described below in the subsections.
4.1 End-to-end-trained Deep Learning Models
We initialized all three models that were trained in an end-to-end fashion (i.e., EfficientNetV2, ViT and Swin Transformer) with ImageNet-pretrained weights, and the original classification head was replaced with a custom binary classifier consisting of a single linear layer. Specifically, we used the vit_large_patch16_224 model from the timm library, a ViT with 224×224 input resolution, 16×16 patch size, and 86.6 million parameters. The Swin Transformer variant used was swin_base_patch4_window7_224 which features a 4×4 patch size and 7×7 shifted attention windows, comprising 87.8 million parameters. For EfficientNetV2, we employed efficientnetv2_m from timm, which has approximately 54.1 million parameters. This protocol was applied identically to all three models to enable fair comparison of their performance on the atypical mitosis classification task.
@@ 4.2 Linear Probing with Foundation Models
We conducted comprehensive evaluation of seven foundation models using a standardized linear probing method on the AMi-Br dataset for atypical versus normal mitosis classification. For each model, we extracted high-dimensional feature embeddings from histopathology images and trained a single linear classification layer while keeping the pre-trained feature extractor frozen. This method was consistently applied across all six foundation models to ensure fair comparison of their feature representation capabilities on the atypical classification task.
@@ 4.3 LoRA Fine-tuning of Foundation Models
Table 3: Performance of different models across three test datasets. Values are reported as mean $\pm$ standard deviation.
Model AMi-Br AtNorM-Br AtNorM-MD Balanced Acc. AUROC Balanced Acc. AUROC Balanced Acc. AUROC EfficientNetV2 $0.7474\pm 0.0167$ $0.8315\pm 0.0168$ $0.7283\pm 0.0168$ $0.8110\pm 0.0084$ $0.6975\pm 0.0277$ $0.7604\pm 0.0217$ Vision Transformer $0.7899\pm 0.0268$ $0.8872\pm 0.0107$ $\textbf{0.7788}\pm 0.0229$ $\textbf{0.8710}\pm 0.0075$ $0.7494\pm 0.0342$ $0.8720\pm 0.0117$ Swin Transformer $0.8052\pm 0.0124$ $\textbf{0.9029}\pm 0.0112$ $0.7752\pm 0.0134$ $0.8683\pm 0.0113$ $\textbf{0.7723}\pm 0.0142$ $\textbf{0.8806}\pm 0.0087$ UNI $0.6339\pm 0.0189$ $0.6949\pm 0.0275$ $0.6406\pm 0.0339$ $0.7028\pm 0.0310$ $0.5510\pm 0.0123$ $0.5892\pm 0.0118$ UNI2-h $0.6220\pm 0.0095$ $0.6765\pm 0.0157$ $0.6482\pm 0.0214$ $0.7059\pm 0.0230$ $0.5773\pm 0.0212$ $0.6119\pm 0.0357$ Virchow $0.6029\pm 0.0132$ $0.6520\pm 0.0168$ $0.5969\pm 0.0155$ $0.6733\pm 0.0266$ $0.5623\pm 0.0235$ $0.6184\pm 0.0566$ Virchow2 $0.6089\pm 0.0192$ $0.6637\pm 0.0214$ $0.6236\pm 0.0296$ $0.6795\pm 0.0344$ $0.5694\pm 0.0173$ $0.6046\pm 0.0324$ Prov-Gigapath $0.6097\pm 0.0177$ $0.6435\pm 0.0188$ $0.5893\pm 0.0074$ $0.6187\pm 0.0193$ $0.5558\pm 0.0125$ $0.6039\pm 0.0244$ H-Optimus-0 $0.6519\pm 0.0089$ $0.7098\pm 0.0062$ $0.6191\pm 0.0156$ $0.7058\pm 0.0140$ $0.5541\pm 0.0323$ $0.5786\pm 0.0458$ H-Optimus-1 $0.6088\pm 0.0120$ $0.6705\pm 0.0094$ $0.6152\pm 0.0305$ $0.6868\pm 0.0195$ $0.6051\pm 0.0142$ $0.6596\pm 0.0294$ UNI (LoRA) $0.7952\pm 0.0092$ $0.8839\pm 0.0059$ $0.7183\pm 0.0226$ $0.7979\pm 0.0142$ $0.7069\pm 0.0278$ $0.8222\pm 0.0224$ UNI2-h (LoRA) $0.7138\pm 0.0121$ $0.8153\pm 0.0211$ $0.7301\pm 0.0152$ $0.8228\pm 0.0143$ $0.6914\pm 0.0321$ $0.7616\pm 0.0415$ Virchow (LoRA) $0.7878\pm 0.0250$ $0.8891\pm 0.0150$ $0.7696\pm 0.0198$ $0.8540\pm 0.0200$ $0.7705\pm 0.0287$ $0.8641\pm 0.0247$ Virchow2 (LoRA) $\mathbf{0.8135}\pm 0.0145$ $0.9026\pm 0.0051$ $0.7632\pm 0.0190$ $0.8579\pm 0.0117$ $0.7424\pm 0.0305$ $0.8503\pm 0.0171$ Prov-Gigapath (LoRA) $0.7602\pm 0.0113$ $0.8682\pm 0.0122$ $0.7263\pm 0.0296$ $0.8077\pm 0.0184$ $0.7007\pm 0.0228$ $0.8073\pm 0.0259$ H-Optimus-0 (LoRA) $0.7846\pm 0.0169$ $0.8721\pm 0.0148$ $0.7044\pm 0.0383$ $0.7921\pm 0.0271$ $0.6549\pm 0.0281$ $0.7879\pm 0.0163$ H-Optimus-1 (LoRA) $0.7762\pm 0.0110$ $0.8611\pm 0.0083$ $0.7658\pm 0.0298$ $0.8524\pm 0.0100$ $0.7396\pm 0.0202$ $0.8389\pm 0.0104$

### v3

@@ Abstract
Atypical mitosis marks a deviation in the cell division process that has been shown be an independent prognostic marker for tumor malignancy. However, atypical mitosis classification remains challenging due to low prevalence, at times subtle morphological differences from normal mitotic figures, low inter-rater agreement among pathologists, and class imbalance in datasets. Building on the Atypical Mitosis dataset for Breast Cancer (AMi-Br), this study presents a comprehensive benchmark comparing deep learning approaches for automated atypical mitotic figure (AMF) classification, including end-to-end fine-tuned deep learning models, foundation models with linear probing, and foundation models fine-tuned with low-rank adaptation (LoRA). For rigorous evaluation, we further introduce two new held-out AMF datasets - AtNorM-Br, a dataset of mitotic figures from the TCGA breast cancer cohort, and AtNorM-MD, a multi-domain dataset of mitotic figures from a subset of the MIDOG++ training set. We found average balanced accuracy values of up to 0.8135, 0.7788, and 0.7723 on the in-domain AMi-Br and the out-of-domain AtNorm-Br and AtNorM-MD datasets, respectively. Our work shows that atypical mitotic figure classification, while being a challenging problem, can be effectively addressed through the use of recent advances in transfer learning and model fine-tuning techniques. We make all code and data used in this paper available in this github repository: https://github.com/DeepMicroscopy/AMi-Br_Benchmark.
@@ 1 Introduction
Mitosis is the process where a cell replicates its genetic material (DNA) and then divides into two identical daughter cells. This process enables the proliferation of cells, supporting physiological tissue growth and facilitating the replacement of old, damaged, or worn-out cells. The dividing cell may be observed microscopically as a mitotic figure (MF) with specific morphologies resembling the highly regulated cell division phases. However, cell proliferation does not occur only under physiological conditions; it is also a hallmark of cancer, where cell division is dysregulated and increased. Subsequently, quantifying cell proliferation is essential for estimating the aggressiveness of many human and animal cancers, including human breast cancer (Fitzgibbons and Connolly, 2023; Kiupel et al., 2011; Louis et al., 2016). The routine method in the assessment of the growth rate is the mitotic count, which reflects the number of MFs within a specific tumor area, selected from a whole slide image (WSI). Elevated mitotic counts are associated with an increased risk of distant metastasis and higher risk of death (Medri et al., 2003).
@@
At the time of this writing, there have been very few studies investigating deep learning for AMFs (Aubreville et al., 2023a; Fick et al., 2024; Bertram et al., 2025). The automated classification of atypical vs. normal MFs is complicated by the substantial variation of morphologies, often with some mitotic figures displaying overlapping features of both normal and AMFs (Donovan et al., 2021; Bertram et al., 2023), and the low frequency of AMF, resulting in class imbalance. Given the prognostic relevance of differentiating normal vs. atypical MFs, standardized benchmarks, such as publicly available datasets, evaluation protocols, and baseline models, may serve as facilitators to push further developments in this field that tackle these challenges. In this context, foundation models offer a promising direction. These models, typically vision transformer (ViT)-based and pre-trained with self-supervised learning on millions of histopathology tiles, have shown strong performance across various medical imaging tasks, including classification (Campanella et al., 2025; Breen et al., 2025). When subsequently fine-tuned on comparatively small, task-specific labelled datasets, they are particularly effective in small data regimes. These models are typically employed either via linear probing, where a lightweight classifier is trained on top of a frozen backbone, or through parameter-efficient fine-tuning techniques like Low Rank Adaptation (LoRA) (Hu et al., 2022). However, comprehensive studies leveraging these techniques for atypical vs. normal mitosis classification are still lacking.
The contribution of this work is two-fold: First, we present three datasets, designed for the automated identification of AMFs to establish suitable datasets that reflect the unique challenges of this task. Second, we provide a comprehensive benchmark comparing end-to-end fine-tuned deep learning architectures, foundation models with linear probing, and foundation models fine-tuned with LoRA for the task of atypical versus normal mitotic figure classification.
@@ 2 Datasets
Being the largest of the three datasets used in this study, we use data from the AMi-Br dataset for training in all evaluations. This allows us to use the other two newly introduced datasets as independent held-out sets. It has been shown that data distribution shifts (domain shifts) impede the performance of deep learning models considerably (Stacke et al., 2020; Aubreville et al., 2021). With these datasets, we expect varying degrees of domain shifts, and which allows us to investigate generalization performance. Only a part of the AMi-Br dataset is used for training the models for binary classification of atypical vs. normal mitoses, and the rest is used as another held-out test dataset. The dataset statistics, including total number of atypical and normal MFs in each, annotation type, source dataset(s), species involved etc. have been summarized in Table 1.
@@
Property AMi-Br AtNorM-Br AtNorM-MD
Total MFs 3,720 746 2,107
Atypical 832 128 219
Normal 2,888 618 1,888
AMF Rate (%) 22.4% 17.2% 10.4%
Annotation Type 3-expert vote Single expert 5-expert vote
@@
Source Dataset(s) TUPAC16, MIDOG21 TCGA (BRCA) MIDOG++
Species Human Human Human + Canine
@@ 2.1 AMi-Br
2.2 AtNorM-Br
The third dataset, Atypical and Normal Mitosis (AtNorM)-Br, also made available publicly within the scope of this work, contains 746 MF instances from 179 patients from the breast cancer (BRCA) cohort of the The Cancer Genome Atlas (TCGA) (Lingle et al., 2016). TCGA contains images from various sources and with partially mixed quality, stemming from differences in staining protocols, scanning equipment, and tissue preparation across institutions, and is thus also a valuable asset for the assessment of generalization. This dataset was annotated by a single expert with high experience in AMF classification, according to which 128 MFs were found to be atypical and the remaining 618 were classified as normal, yielding an AMF rate of 17.16%.
@@ 2.3 AtNorM-MD
The second dataset, AtNorM-MultiDomain (MD) extends beyond human breast cancer and covers six domains, spanning both human and canine tumors. These include canine lung cancer, canine lymphoma, canine cutaneous mast cell tumor, human neuroendocrine tumor, canine soft tissue sarcoma, and human melanoma. It was sampled randomly from all but the human breast cancer domains of the publicly available MIDOG++ dataset (Aubreville et al., 2023c). It comprises 2,107 MFs from 70 patients, with 219 (10.4%) being atypical. Labeling was performed as a majority vote of five pathology experts, three of whom were board-certified. We found total agreement per object by all five experts in 1,466 (69.6%) of the cases.
@@ 3 Methods
3.1 End-to-end fine-tuned Baseline Deep Learning Models
To establish robust baselines for atypical vs. normal mitoses classification, we evaluate three widely-used deep learning architectures – EfficientNetV2 (Tan and Le, 2021), ViT (Dosovitskiy et al., 2020), and Swin Transformer (Liu et al., 2021). All of these models have demonstrated strong performance across various medical imaging datasets and tasks. EfficientNetV2 is a well-established convolutional neural network (CNN) known for its strong performance and efficiency in various classification tasks. It serves as a representative of traditional CNN-based approaches. In contrast, the ViT represents a major shift from convolutional architectures to transformer-based models. A ViT processes images as sequences of patches and uses self-attention mechanisms to model global relationships across the input. This global receptive field allows the ViT to capture long-range dependencies, which can be particularly useful for complex histopathological patterns. The Swin Transformer further advances the transformer-based approach by introducing a hierarchical architecture with shifted windows. This architecture enables the model to compute attention locally while gradually building up to global representations, providing a balance between computational efficiency and the ability to capture both local details and global structure. Swin Transformers have been particularly effective in dense prediction tasks and high-resolution image analysis, making it well-suited for mitotic figure classification (Liu et al., 2021). The selected model architectures were chosen to be of similar size in parameters. Each model was fine-tuned in an end-to-end manner and evaluated on three held-out test datasets: a random, patient-stratified split of the AMi-Br dataset, accounting for $\approx$ 22 % of the samples in the dataset (samples not used in training or validation), and the entire AtNorM-MD and AtNorM-Br datasets.
@@ 3.2 Foundation Models
Foundation models are models that are trained on large and diverse corpora of data, often using unsupervised techniques like self supervision. The goal of foundation models is to train feature extractors that generalize well and are easily adaptable to down-stream tasks. In the field of computational pathology, we have seen a very strong incline towards publicly available foundation models, calling for an investigation on the task of AMF classification. We compare eight state-of-the-art models – UNI (Chen et al., 2024), UNI2-h (Chen et al., 2024; MahmoodLab, 2025), Virchow (Vorontsov et al., 2024), Virchow2 (Zimmermann et al., 2024), Prov-Gigapath (Xu et al., 2024), H-Optimus-0 (Saillard et al., 2024), H-Optimus-1 (Bioptimus, 2025), and H0-mini (Filiot et al., 2025). An overview of selected details for each model such as training dataset size, size in parameters and model type for training is provided in Table 2. All selected foundation models are ViT-based and have been pre-trained using the DINOv2 algorithm (Oquab et al., 2023) on datasets comprising of hundreds of thousands to multiple millions of WSIs.
@@ 4 Experimental Setup
We set up detailed experiments for the above cases. All training was performed exclusively on the AMi-Br dataset. The image preprocessing for the binary classification experiments included resizing all patches, originally of the size $128\times 128$, to $224\times 224$ pixels. Training augmentations consisted of random horizontal flip, rotation, color jitter, and random resized crop, followed by normalization. For models initialized from ImageNet-pretrained weights and fine-tuned end-to-end, we used the standard ImageNet mean and standard deviation for normalization. For the computational pathology foundation models used in the linear probing and LoRA experiments, we instead followed the pre-processing recommended for each model, i.e., normalizing with the model-specific mean and standard deviation used during their pre-training, rather than assuming ImageNet statistics. Validation images were only resized and normalized. The best checkpoint for each fold was selected based on validation balanced accuracy. The balanced accuracy for binary classification is the average of sensitivity and specificity and can be mathematically stated as follows:
$\text{Balanced Accuracy}=\frac{\text{Sensitivity}+\text{Specificity}}{2}$
The balanced accuracy used as the primary metric because of the class imbalance between the atypical and normal MFs in the dataset, with a fixed decision threshold of 0.5 on the predicted probabilities. All experiments were performed using 5-fold cross validation, where stratification of the data was based on the slide (patient) to ensure that all patches from a given slide appeared in only one split (train or validation), thus preventing data leakage. To address class imbalance during training, we used a weighted random sampler that assigns higher sampling weights to under-represented classes and lower weights to frequent classes, with each sample’s weight set inversely to its class frequency. We used weighted random sampling because it was proven to give the best results in previous work (Bertram et al., 2025). All models were trained using the Adam optimizer with L2 regularization (learning rate: $5\times 10^{-5}$, weight decay: $1\times 10^{-5}$) and cross-entropy loss. We applied early stopping with a patience of 15 epochs based on validation balanced accuracy, and used learning rate scheduling on plateau (factor: 0.5, patience: 3, minimum LR: $1\times 10^{-7}$). Each model was trained for up to 100 epochs with a batch size of 8. The rest of the case-specific details are described below in the subsections.
4.1 End-to-end fine-tuned Baseline Deep Learning Models
We initialized all three models (i.e., EfficientNetV2, ViT and Swin Transformer) that were fine-tuned in an end-to-end fashion, essentially doing a full finetuning of all the layers with ImageNet-pretrained weights, and the original classification head was replaced with a custom binary classifier consisting of a single linear layer. Specifically, we used the vit_large_patch16_224 model from the timm library (Wightman, 2019), a ViT with 224×224 input resolution, 16×16 patch size, and 86.6 million parameters. The Swin Transformer variant used was swin_base_patch4_window7_224 which features a 4×4 patch size and 7×7 shifted attention windows, comprising 87.8 million parameters. For EfficientNetV2, we employed efficientnetv2_m from timm, which has approximately 54.1 million parameters. This protocol was applied identically to all three models to enable fair comparison of their performance on the atypical mitosis classification task.
@@ 4.2 Linear Probing with Foundation Models
We conducted comprehensive evaluation of seven foundation models using a standardized linear probing method for atypical versus normal mitosis classification. For each model, we extracted high-dimensional feature embeddings from histopathology images and trained a single linear classification layer while keeping the pre-trained feature extractor frozen. This method was consistently applied across all six foundation models to ensure fair comparison of their feature representation capabilities on the atypical classification task.
@@ 4.3 LoRA Fine-tuning of Foundation Models
Table 3: Classification performance of the evaluated models on the AMi-Br, AtNorM-Br, and AtNorM-MD test sets. For each dataset, balanced accuracy and AUROC are reported as mean $\pm$ standard deviation across cross-validation folds. All models were trained exclusively on the AMi-Br dataset, while evaluation was performed on the respective test sets. Within each column, bold values denote the best-performing model for that dataset and in the respective group of models.
Model AMi-Br AtNorM-Br AtNorM-MD Balanced Acc. AUROC Balanced Acc. AUROC Balanced Acc. AUROC EfficientNetV2 (Baseline) $0.7474\pm 0.0167$ $0.8315\pm 0.0168$ $0.7283\pm 0.0168$ $0.8110\pm 0.0084$ $0.6975\pm 0.0277$ $0.7604\pm 0.0217$ Vision Transformer (Baseline) $0.7899\pm 0.0268$ $0.8872\pm 0.0107$ $\textbf{0.7788}\pm 0.0229$ $\textbf{0.8710}\pm 0.0075$ $0.7494\pm 0.0342$ $0.8720\pm 0.0117$ Swin Transformer (Baseline) $\textbf{0.8052}\pm 0.0124$ $\textbf{0.9029}\pm 0.0112$ $0.7752\pm 0.0134$ $0.8683\pm 0.0113$ $\textbf{0.7723}\pm 0.0142$ $\textbf{0.8806}\pm 0.0087$ UNI $\textbf{0.6339}\pm 0.0189$ $\textbf{0.6949}\pm 0.0275$ $0.6406\pm 0.0339$ $0.7028\pm 0.0310$ $0.5510\pm 0.0123$ $0.5892\pm 0.0118$ UNI2-h $0.6220\pm 0.0095$ $0.6765\pm 0.0157$ $\textbf{0.6482}\pm 0.0214$ $0.7059\pm 0.0230$ $0.5773\pm 0.0212$ $0.6119\pm 0.0357$ Virchow $0.6029\pm 0.0132$ $0.6520\pm 0.0168$ $0.5969\pm 0.0155$ $0.6733\pm 0.0266$ $0.5623\pm 0.0235$ $0.6184\pm 0.0566$ Virchow2 $0.6089\pm 0.0192$ $0.6637\pm 0.0214$ $0.6236\pm 0.0296$ $0.6795\pm 0.0344$ $0.5694\pm 0.0173$ $0.6046\pm 0.0324$ Prov-Gigapath $0.6097\pm 0.0177$ $0.6435\pm 0.0188$ $0.5893\pm 0.0074$ $0.6187\pm 0.0193$ $0.5558\pm 0.0125$ $0.6039\pm 0.0244$ H-Optimus-0 $0.6319\pm 0.0153$ $0.6801\pm 0.0181$ $0.6369\pm 0.0147$ $\textbf{0.7090}\pm 0.0163$ $0.5821\pm 0.0195$ $0.6142\pm 0.0207$ H-Optimus-1 $0.5943\pm 0.0132$ $0.6518\pm 0.0152$ $0.6271\pm 0.0168$ $0.6777\pm 0.0275$ $\textbf{0.5966}\pm 0.0317$ $\textbf{0.6462}\pm 0.0461$ UNI (LoRA) $0.7952\pm 0.0092$ $0.8839\pm 0.0059$ $0.7183\pm 0.0226$ $0.7979\pm 0.0142$ $0.7069\pm 0.0278$ $0.8222\pm 0.0224$ UNI2-h (LoRA) $0.7138\pm 0.0121$ $0.8153\pm 0.0211$ $0.7301\pm 0.0152$ $0.8228\pm 0.0143$ $0.6914\pm 0.0321$ $0.7616\pm 0.0415$ Virchow (LoRA) $0.7878\pm 0.0250$ $0.8891\pm 0.0150$ $\textbf{0.7696}\pm 0.0198$ $0.8540\pm 0.0200$ $\textbf{0.7705}\pm 0.0287$ $\textbf{0.8641}\pm 0.0247$ Virchow2 (LoRA) $\mathbf{0.8135}\pm 0.0145$ $\textbf{0.9026}\pm 0.0051$ $0.7632\pm 0.0190$ $\textbf{0.8579}\pm 0.0117$ $0.7424\pm 0.0305$ $0.8503\pm 0.0171$ Prov-Gigapath (LoRA) $0.7602\pm 0.0113$ $0.8682\pm 0.0122$ $0.7263\pm 0.0296$ $0.8077\pm 0.0184$ $0.7007\pm 0.0228$ $0.8073\pm 0.0259$ H-Optimus-0 (LoRA) $0.7846\pm 0.0169$ $0.8721\pm 0.0148$ $0.7044\pm 0.0383$ $0.7921\pm 0.0271$ $0.6549\pm 0.0281$ $0.7879\pm 0.0163$ H-Optimus-1 (LoRA) $0.7762\pm 0.0110$ $0.8611\pm 0.0083$ $0.7658\pm 0.0298$ $0.8524\pm 0.0100$ $0.7396\pm 0.0202$ $0.8389\pm 0.0104$
Table 4: p-values for group comparisons of balanced accuracies within each dataset.
Dataset Linear Probing vs LoRA (Wilcoxon signed-rank) Baseline vs Linear Probing (Mann–Whitney U) Baseline vs LoRA (Mann–Whitney U)
AMi-Br $5.82\text{\times}{10}^{-11}$ $2.91\text{\times}{10}^{-08}$ $0.73$
AtNorM-Br $1.16\text{\times}{10}^{-10}$ $2.91\text{\times}{10}^{-08}$ $0.06$
AtNorM-MD $5.82\text{\times}{10}^{-11}$ $2.91\text{\times}{10}^{-08}$ $0.09$